The patient is to receive fentanyl 25 mg IV. Available is fentanyl 50 mg/ml. How much time is required to administer the medication?
Answer #1

Ans) Ordered dose: Fentanyl 25 mg IV

Available: Fentanyl: 50 mg/ml

So 50mg = 1 ml

Then 25 mg= x

Let x be the time required to administer the ordered dose

x= 25/50 = 1/2 hrs = 30 mins.

The time required to administer the dose is 30 mins.